Five more than four times a number is 53
Answer by jim_thompson5910(35256)

"Five more than four times a number is 53"

translates to

4x + 5 = 53

where x is the unknown number

Solve for x...

4x+5=53

4x=53-5

4x=48

x=(48)/(4)

x=12

So the unknown number is 12